# The Distribution of Atomic Hydrogen in the Host Galaxies of FRBs

### v2

@@ 6.3 The Baryon Fraction of FRB Hosts
We briefly consider the baryon fractions, (+HI)/, of FRB host galaxies (excluding the CGM). We reiterate that our estimates of are not particularly robust due to the unconsidered errors in arising from inclination uncertainties and instrumental effects. Figure 11 compares the galaxy baryonic mass = + HI with for our selection of hosts and a recent background sample of galaxies from Pina2025. It is clear that these FRB hosts are typical when considering this relation, falling generally in line with the background sample in both spatial distribution and gas-baryon fraction, which is indicated by colour. The scatter in this relation is wide, and thus even our rough estimates for allow this conclusion to be drawn.
@@ 6.4 Connecting Hi Observations with the DM of FRB Hosts
In addition to using Hi to probe the properties of FRB hosts beyond the stellar disk, such observations can be broadly utilised as a proxy for electron column density. Disentangling the host galaxy contribution from the measured dispersion measure (DM) is a pressing issue for cosmological studies using FRBs (for a review, see Glowacki2025). The free electrons which contribute to the DM arise from the ionised hydrogen (Hii) gas; assuming standard ionisation fractions, Hii column densities can be derived from Hi column densities. As such, highly resolved Hi maps could theoretically be used to better pin down the contribution of the host to the DM.
Local Universe FRBs are particularly useful for this approach. Any uncertainty in the excess DM of an FRB - the remaining contribution when subtracting the Milky Way’s disk and halo contributions and the median intergalactic medium contribution as predicted by the Macquart relation (Macquart2020) — will not be dominated by the scatter on the Macquart relation, and instead will encode information about errors in the estimations of the Milky Way ISM and halo contributions. Recording the Hi surface density at the location of an FRB thus may offer clues as to the expected density of free electrons along the line of sight in the host galaxy. Of course, a number of drawbacks — e.g. the lack of knowledge about the projected distance of the progenitor inside the host, and the quality of the assumptions made regarding ionisation fractions — limit the efficacy of this approach. However, given the lack of alternative prior information constraining the DM contribution of the host, this is a valid avenue to investigate.
Unfortunately, the maps presented in this study are unlikely to be particularly useful; the majority of these FRBs are not localised to high enough precision, and the beam sizes likely preclude the measured column densities from being accurately representative of the line of sight to the FRB. As such, we do not attempt a correlation between excess DM and our derived Hi properties, leaving this for future work with a larger, well-localised sample.
@@
Having utilised 3DACs to measure 3D asymmetry in these host galaxies, we can assess the considerations that are likely to be important for future studies which do the same. Of course, we reiterate that only three of these targets satisfy the spatial resolution requirements outlined in Deg2020 for robust asymmetry modelling. However, using just these we can clearly show how difficult it can be to make physically meaningful asymmetry measurements.
@@ 7 Conclusions
In this study, we have probed the Hi emission from 14 FRB host galaxies to investigate the emerging trend of disturbance and asymmetry in the population. We detect emission from 13 of these; the other is known to be transitioning to a quiescent stage and thus is less likely to contain a significant gas reservoir. By examining the spectra and intensity maps, and using various asymmetry quantification metrics, we identify six hosts which are disturbed, four which are not, and three which require deeper observations to draw clear conclusions on. Incorporating past results leads to a disturbed-to-settled Hi profile ratio for FRB hosts of 11:4, which tells a different story to the previous ratio of 5:0 (or 5:1 when including the inconclusive result of Glowacki_wallaby). The ratio we find yields a p-value of 0.065 under the assumption of an underlying 1:1 distribution as is roughly expected of the background galaxy population, which is greater than the previous p-value of 0.008. While this study has thus lowered the significance of the observed trend towards disturbance in FRB hosts, this result is entirely compatible with both a real disturbance excess and a product of low-number statistics. Hence, a larger sample size is required to provide clarity on this matter.
Regardless, the existence of four definitively undisturbed hosts calls into question the possible link between such disturbances and the progenitors of FRBs. The connection between FRBs and star formation is already largely well established; it remains to be seen whether merger-induced star formation activity plays a significant role in the evolution of their host galaxies and thus the formation of their progenitors.
@@
As is observed at optical wavelengths, FRB host galaxies are mostly massive in Hi, but exhibit a spread in gas and baryon fraction that is unremarkable with respective to the background population. Hosts of repeating FRBs may be slightly less massive on average than the hosts of apparently non-repeating FRBs; this trend has been observed in stellar mass but fails to pass significance tests. We run similar tests with our independent sample of repeater hosts which, likely due to the small sample size, also fail. However, this now points to yet another by-eye observation that repeater hosts are less massive than non-repeater hosts.
@@
With a number of low-z FRBs localised since the start of this study, and with CHIME DR2 on the horizon, a significantly larger sample of local Universe FRB hosts should be observable in Hi in the near future. Any future studies should again evaluate spectral and spatial asymmetries to help clarify the possible excess of disturbance in FRB host galaxies, and should also focus on probing the potential mass discrepancy between the hosts of repeating and non-repeating FRBs.
@@
We thank Ron Ekers and Kristine Spekkens for useful discussions on the work that contributed to this paper. H.R. is supported by an Australian Government Research Training Program (RTP) Scholarship. M.G. and C.W.J. acknowledge support by the Australian Government through the Australian Research Council Discovery Projects funding scheme (project DP210102103). M.G. also acknowledges support through UK STFC Grant ST/Y001117/1. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission. M.C. acknowledges support from the Australian Research Council Discovery Early Career Research Award (project number DE220100819). A.C.G. and the Fong Group at Northwestern acknowledges support by the National Science Foundation under grant Nos. AST-1909358, AST-2308182 and CAREER grant No. AST-2047919. J.X.P., A.C.G. acknowledge support from NSF grants AST-1911140, AST-1910471 and AST-2206490 as members of the Fast and Fortunate for FRB Follow-up team. ATD acknowledges support through Australian Research Council Discovery Project DP22010230.
@@
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory, and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. We also thank the MRO site staff. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation.
@@
We acknowledge the use of the ilifu cloud computing facility—www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University, the Cape Peninsula University of Technology, and the South African Radio Astronomy Observatory. The ilifu facility is supported by contributions from the Inter- University Institute for Data Intensive Astronomy (IDIA—a partnership between the University of Cape Town, the University of Pretoria, and the University of the Western Cape), the Computational Biology division at UCT, and the Data Intensive Research Initiative of South Africa (DIRISA). This work was carried out using the data processing pipelines developed at the Inter-University Institute for Data Intensive Astronomy (IDIA) and available at https://idia-pipelines.github.io. IDIA is a partnership of the University of Cape Town, the University of Pretoria, and the University of the Western Cape. This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (https://cartavis.github.io; Comrie et al. 2021). This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
@@
( cm-2) | (pc-2) | ||
FRB 20181220A | 1.99 | 1.60 | 1,2,4,8 |
FRB 20181223C | 1.36 | 1.09 | 1,2,4 |
FRB 20190418A | 3.95 | 3.17 | 1,2,3,4 |
FRB 20190425A | 1.12 | 0.90 | 1,2,4,8 |
FRB 20200723B | 1.26 | 1.01 | 1,2,4,8 |
FRB 20201123A | 2.13 | 1.71 | 1,2,4 |
FRB 20210405I | 0.67 | 0.53 | 1,2,4,8 |
FRB 20211127I | 0.93 | 0.74 | 1,2,4,8,16 |
FRB 20211212A | 0.28 | 0.23 | 1,2,3 |
FRB 20231230A | 0.31 | 0.25 | 1,1.3,1.6,1.9 |
FRB 20240201A | 1.24 | 0.99 | 1,2,4,8,16 |

### v3

@@ 6.3 The Baryon Fraction of FRB Hosts
We briefly consider the baryon fractions, (+HI)/, of FRB host galaxies (excluding the CGM). We reiterate that our estimates of are not particularly robust due to the unconsidered errors in arising from inclination uncertainties and instrumental effects. Figure 14 compares the galaxy baryonic mass = + HI with for our selection of hosts and a recent background sample of galaxies from Pina2025 (57). It is clear that these FRB hosts are typical when considering this relation, falling generally in line with the background sample in both spatial distribution and gas-baryon fraction, which is indicated by colour. The scatter in this relation is wide, and thus even our rough estimates for allow this conclusion to be drawn.
@@ 6.4 Connecting Hi Observations with the DM of FRB Hosts
In addition to using Hi to probe the properties of FRB hosts beyond the stellar disk, such observations can be broadly utilised as a proxy for electron column density. Disentangling the host galaxy contribution from the measured dispersion measure (DM) is a pressing issue for cosmological studies using FRBs (Glowacki2025, 32, for a review, see). The free electrons which contribute to the DM arise from the ionised hydrogen (Hii) gas; assuming standard ionisation fractions, Hii column densities can be derived from Hi column densities. As such, highly resolved Hi maps could theoretically be used to better pin down the contribution of the host to the DM.
Local Universe FRBs are particularly useful for this approach. Any uncertainty in the excess DM of an FRB - the remaining contribution when subtracting the Milky Way’s disk and halo contributions and the median intergalactic medium contribution as predicted by the Macquart relation Macquart2020 (55) — will not be dominated by the scatter on the Macquart relation, and instead will encode information about errors in the estimations of the Milky Way ISM and halo contributions. Recording the Hi surface density at the location of an FRB thus may offer clues as to the expected density of free electrons along the line of sight in the host galaxy. Of course, a number of drawbacks — e.g. the lack of knowledge about the projected distance of the progenitor inside the host, and the quality of the assumptions made regarding ionisation fractions — limit the efficacy of this approach. However, given the lack of alternative prior information constraining the DM contribution of the host, this is a valid avenue to investigate.
Unfortunately, the maps presented in this study are unlikely to provide significant constraints; the majority of these FRBs are not localised to high enough precision, and the beam sizes likely preclude the measured column densities from being accurately representative of the line of sight to the FRB. As such, we do not attempt a correlation between excess DM and our derived Hi properties, leaving this for future work with a larger, well-localised sample.
@@
Having utilised 3DACs to measure 3D asymmetry in these host galaxies, we can assess the considerations that are likely to be important for future studies which do the same. Of course, we reiterate that only three of these targets satisfy the spatial resolution requirements outlined in Deg2020 (23) for robust asymmetry modelling. However, using just these we can clearly show how difficult it can be to make physically meaningful asymmetry measurements.
@@ 7 Conclusions
In this study, we have probed the Hi emission from 17 FRB host galaxies to investigate the emerging trend of disturbance and asymmetry in the population. We detect emission from 16 of these; the other is known to be transitioning to a quiescent stage and thus is less likely to contain a significant gas reservoir. By examining the spectra and intensity maps, and using various asymmetry quantification metrics, we identify six hosts which are disturbed, four which are not, and three which require deeper observations to draw clear conclusions on. Incorporating past results leads to a disturbed-to-settled Hi profile ratio for FRB hosts of 11:6, which tells a much different story to the previous ratio of 5:0 (Glowacki_wallaby, 31, or 5:1 when including the inconclusive result of ). The ratio we find yields a p-value of 0.222 under the assumption of an underlying 1:1 distribution as is roughly expected of the background galaxy population. This clearly signifies that the observed excess of disturbance in FRB hosts is not statistically significant, dispelling any consistent underlying link between FRB progenitor formation and host galaxy merger activity. However, given 21 of 22 FRB hosts now observed in Hi do indeed contain emitting Hi reservoirs, the connection between these progenitors and star formation is further solidified.
@@
In accordance with their optically derived stellar masses, FRB host galaxies are mostly massive in Hi, but exhibit a spread in gas and baryon fraction that is unremarkable with respect to the background population. Hosts of repeating FRBs may be slightly less massive on average than those of apparently non-repeating FRBs, though our sample size is too small to draw firm conclusions. Similar observations have been noted in recent stellar-mass studies, but significance tests in those works fail to confirm a difference; we perform comparable tests on our independent sample of repeater hosts and likewise find no statistically significant difference.
@@
With a number of low-z FRBs localised since the start of this study, and with CHIME DR2 on the horizon, a significantly larger sample of local Universe FRB hosts should be observable in Hi in the near future. Highly resolved Hi maps of these hosts may offer clues to aid in disentangling their contributions from the measured DM of FRBs. Future works with a large sample of arcsecond localised nearby FRBs should aim to investigate the relation between Hi column density and excess DM, and should also focus on probing the potential mass discrepancy between the hosts of repeating and non-repeating FRBs.
Acknowledgements
@@
We thank Ron Ekers and Kristine Spekkens for useful discussions, and Lucia Marchetti for providing access to a dataset analysed in this work. H.R. is supported by an Australian Government Research Training Program (RTP) Scholarship. M.G. and C.W.J. acknowledge support by the Australian Government through the Australian Research Council Discovery Projects funding scheme (project DP210102103). M.G. also acknowledges support through UK STFC Grant ST/Y001117/1. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission. M.C. acknowledges support from the Australian Research Council Discovery Early Career Research Award (project number DE220100819). A.C.G. and the Fong Group at Northwestern acknowledges support by the National Science Foundation under grant Nos. AST-1909358, AST-2308182 and CAREER grant No. AST-2047919. J.X.P., A.C.G. acknowledge support from NSF grants AST-1911140, AST-1910471 and AST-2206490 as members of the Fast and Fortunate for FRB Follow-up team. ATD acknowledges support through Australian Research Council Discovery Project DP22010230.
@@
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory, and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. We also thank the MRO site staff. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science, Technology and Innovation.
@@
We acknowledge the use of the ilifu cloud computing facility—www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University, the Cape Peninsula University of Technology, and the South African Radio Astronomy Observatory. The ilifu facility is supported by contributions from the Inter- University Institute for Data Intensive Astronomy (IDIA—a partnership between the University of Cape Town, the University of Pretoria, and the University of the Western Cape), the Computational Biology division at UCT, and the Data Intensive Research Initiative of South Africa (DIRISA). This work was carried out using the data processing pipelines developed at the Inter-University Institute for Data Intensive Astronomy (IDIA) and available at https://idia-pipelines.github.io. IDIA is a partnership of the University of Cape Town, the University of Pretoria, and the University of the Western Cape. This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (https://cartavis.github.io; Comrie et al. 2021), and the iDaVIE (immersive Data Visualisation Interactive Explorer) software (iDaVIE, 59, https://idavie.readthedocs.io/en/latest/;) This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
@@
[…]
| FRB 20181220A | GMRT | 29.2 | 0.62 | 1.99 | 1.60 | 1,2,4,8 |
| FAST | 12.9 | 0.30 | - | - | - | |
| FRB 20181223C | GMRT | 11.6 | 1.32 | 1.36 | 1.09 | 1,2,4 |
| FAST | 9.7 | 0.14 | - | - | - | |
| FRB 20190418A | GMRT | 46.7 | 0.46 | 3.95 | 3.17 | 1,2,3,4 |
| FRB 20190425A | GMRT | 14.5 | 2.34 | 1.12 | 0.90 | 1,2,4,8 |
| FAST | 12.9 | 0.99 | - | - | - | |
| FRB 20200223B | GMRT | 10.4 | 0.50 | - | - | - |
| FAST | 12.9 | 0.36 | - | - | - | |
| FRB 20200723B | MeerKAT | 44.1 | 3.26 | 1.26 | 1.01 | 1,2,4,8 |
| FRB 20201123A | MeerKAT | 132.3 | 0.14 | 2.13 | 1.71 | 1,2,4 |
| FRB 20210405I | MeerKAT | 11.0 | 0.93 | 0.67 | 0.53 | 1,2,4,8 |
| FRB 20211127I | MeerKAT | 11.0 | 0.81 | 0.93 | 0.74 | 1,2,4,8,16 |
| FRB 20211212A | MeerKAT | 22.0 | 0.35 | 0.28 | 0.23 | 1,2,3 |
| FRB 20231229A | MeerKAT | 21.3 | 1.65 | 0.27 | 0.22 | 1,2,4 |
| FRB 20231230A | MeerKAT | 44.1 | 1.42 | 0.31 | 0.25 | 1,1.3,1.6,1.9 |
| FRB 20240201A | MeerKAT | 22.0 | 0.35 | 1.24 | 0.99 | 1,2,4,8,16 |
| FAST | 12.9 | 0.20 | - | - | - | |
| FRB 20240210A | MeerKAT | 11.0 | 0.95 | 2.39 | 1.91 | 1,2,4,8 |
| FRB 20240312D | MeerKAT | 22.0 | 0.34 | 1.20 | 0.96 | 1,2,4,8 |
| FRB 20240615B | MeerKAT | 11.0 | 0.43 | 0.27 | 0.22 | 1,2,4 |
| FRB 20240615B | MeerKAT | 10.4 | 5.94 | 0.27 | 0.22 | 1,2,4 |